OpenStudy (anonymous):

Consider the following. y=(sqrt8+3x) (a) Write the composite function in the form f(g(x)) by identifying the inner function u = g(x) and the outer function y = f(u). u = g(x) = ? y = f(u) = ? (b) Find the derivative dy/dx. dy/dx = ?

OpenStudy (anonymous):

can you show me step by step please

OpenStudy (anonymous):

is this \[y = \sqrt{8+3x}\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok then inner function \[g(x)=8+3x\]

OpenStudy (anonymous):

yes i got that

OpenStudy (anonymous):

outer function \[f(u)=\sqrt{u}\]

OpenStudy (anonymous):

so what would that be

OpenStudy (anonymous):

so if \[u=8+3x\] you get \[f(g(x))=f(u)=f(8+3x)=\sqrt{8+3x}\]

OpenStudy (anonymous):

yes but whats u?

OpenStudy (anonymous):

\[u=8+3x\] \[\frac{du}{dx}=3\]

OpenStudy (anonymous):

no its \[-22x(5-x^2)^10\]

OpenStudy (anonymous):

\[\frac{dy}{du}=\frac{1}{2\sqrt{u}}\]

OpenStudy (anonymous):

and therefore \[\frac{dy}{dx}=\frac{dy}{du}\frac{du}{dx}=\frac{1}{2\sqrt{u}}\times 3=\frac{3}{2\sqrt{8+3x}}\]

OpenStudy (anonymous):

the original function was \[f(x)=\sqrt{8+3x}\] and \[f'(x)=\frac{3}{2\sqrt{8+3x}}\]

OpenStudy (anonymous):

\[−22x(5−x^2)^{10}\] must be the answer to a different problem.

OpenStudy (anonymous):

it is the answer to "what is the derivative of \[f(x)=(5-x^2)^{11}\]

OpenStudy (anonymous):

ok but ill get back to you when i see what i did